# Pattern formation in a coupled driven diffusive system

### v1

@@ V Discussion and conclusion
In this work, we investigated a model of a driven binary mixture (dubbed the “FLM”) which effectively captures the behavior of the driven Widom-Rowlinson lattice gas (DWRLG) in both the low- and high-density phases, but includes an additional intermediate regime with irregular stripes. In the high-density, ordered phase, earlier studies [1, 2] report standard coarsening dynamics in the absence of a drive, while the addition of a drive leads to the emergence of regular stripes oriented perpendicular to the drive direction. In Sec. II, we briefly characterized the FLM and observed striking similarities with the DWRLG phenomenology, suggesting that both models may belong to the same universality class within a properly chosen parameter space. Notably, we observe that perpendicular stripes emerge even when the system is initialized with parallel stripes [48], indicating a robust stripe-selection mechanism. This reorientation appears to proceed via an instability that is reminiscent of the Mullins-Sekerka instability of the interfaces [48].
@@
Above the critical density, roughly estimated as $\rho_{c}\simeq 0.39$ for $\delta=0$ and $\rho_{\ell}=0.40$ for $\delta=1$, we identify two additional transition densities in the FLM: (i) For $\delta>0$, the density $\rho_{u}$ ($\simeq 0.65$, for $\delta=1$) marks a transition from short-range ordered (irregular) stripes to long-range ordered (regular) stripes [see, e.g., Figs. 1(a,d)- 5(b,c)]. The irregular stripe region, $\rho_{\ell}<\rho<\rho_{u}$, is characterized by pronounced long-wavelength instabilities, which serve as a mechanism to reduce the number of stripes in the system. (ii) For $\delta\geq 0$, $\rho_{g}\simeq 0.90$; marks the onset of a frozen regime for uniform initial conditions [Figs. 1(a,b)], where coarsening arrests and glassy configurations appear. In Sec. II we offer a number of ways to characterize these two regimes.
@@
In the driven low-density regime, the FLM shows indications of a peak in the charge fluctuation (CF) [upper panel in Fig. 6(a)] structure factor at an $L$-independent characteristic wavevector, $q^{*}=0.101\pm 0.001$. The density fluctuation (DF) structure factor [lower panel in Fig. 6(a)] also shows a shoulder at $2q^{*}_{-}$, but remains peaked at $2\pi/L$, consistent with the results reported in Refs. [1, 2]. Examining a snapshot of a typical configuration in this regime [Fig. 6(b)], we see that this microemulsion phase emerges from a tendency to form stripes perpendicular to the drive that is frustrated by the stronger effect of the drive in regions having low particle density.
@@
Deriving a continuum version of the FLM via a gradient and field expansion, we describe the continuum spatiotemporal dynamics of the charge and density fluctuation fields in Sec. III. The analysis yields a system of two (reciprocally) coupled, driven-diffusive, nonlinear partial differential equations, Eq. (III). These equations are consistent with the field equations developed in Ref. [2] for the low-density phase, but include additional terms that are important for describing the high-density regime, namely the terms proportional to $\lambda_{\pm},\widetilde{\lambda}_{\pm}$ and $\Gamma_{\pm}$ in Eq. (III). Using a series of simplifying assumptions, we reduced the full system of equations to a minimal form, given by Eq. (III). The linear coefficients in Eq. (III) are fixed by the GE procedure; Eq. (III), where we see that the charge fluctuation diffusion constant, $D_{-}$, changes sign at $\rho_{c}^{\mathrm{FT}}=0.37\dots$. For comparison, the estimates for the FLM are $\rho_{c}^{(0)}\simeq 0.39$ for the critical density at zero drive ($\delta=0$) and $\rho_{\ell}\simeq 0.40$ as the critical density for (irregular) stripe formation at $\delta=1$.
@@
In Sec. IV, we numerically integrate the field equations. We find that the PSSETD2H method, detailed in the section and in Appendices B and C, is best suited for numerically integrating our particular equations, offering excellent precision and high performance at the cost of reduced flexibility in domain geometry. Moreover, dealiasing techniques are straightforward to implement, effectively minimizing errors from spatial discretization. This is particularly important in the presence of noise  [^18] In fact, we find that it is necessary to use a much finer mesh in space and time than the one we use in Sec. IV to numerically integrate an equation of type (III) with a noise of moderate intensity, $\sigma\approx 10^{-2}$, without recurring to dealiasing techniques.. Interestingly, we found no references in literature that describe these methods in detail, leading us to believe that such a description (and application) could be useful, particularly in the active matter context, since the non-reciprocal Cahn-Hilliard and active model B equations are analogous to Eq. (III). The remainder of Sec. IV provides a qualitative overview of the continuum stochastic field theory, with minimal emphasis on quantitative analysis. A central aspect is that noise plays a crucial role in reproducing the features observed in previous studies. In the low-density regime, we take advantage of the previous perturbative calculations [2] and adopt the previously-derived nonlinear coupling values. Under these conditions, we observe not only a discontinuity in the structure factors $S_{\pm}(\mathbf{q})$ at the origin but also a peak in the charge fluctuation structure factor $S_{-}(\mathbf{q})$ at a nonzero wavevector along the drive direction, an indicator of the microemulsion phase.
@@
Turning to the high-density regime, the undriven ($\delta=0$) case displays standard coarsening dynamics consistent with the LS law in both the lattice models and the field theory. Our main findings for $\delta>0$ are summarized in Figs. 10-11. We identify a regime in the field theory, $g_{0}\sim 0$ and $g_{-}\sim 0.5$, where the equations of motion best reproduce the features of the lattice models (see Fig. 11 (2)). In this regime, as both density and drive increase, the continuum model transitions from irregular to regular stripe configurations, similarly to the FLM. Additionally, in Fig. 13, we uncover several novel regimes not previously reported in the lattice studies, including stripes aligned parallel to the drive and turbulent-like behavior.
@@
A key finding of Sec. IV is that a nonzero difference in the characteristic velocities of the fields, $\Delta v$, is a necessary condition for the formation of stripes aligned perpendicular to the drive. When $\Delta v=0$, such stripes do not form, even when the other drive-dependent couplings $g_{0}$ and $g_{-}$ are nonzero. Moreover, the modulation of the CF field (and, consequently, the DF field) along the drive direction, $q^{*}$, increases with the magnitude of $\Delta v$, in agreement with findings from the low-density regime in [2]; see Fig. 12. However, vertically aligned stripes emerge only when $\Delta v$ exceeds a finite threshold. For instance, at $\rho=0.7$ with fixed nonlinear couplings $g_{-}=0.5$, $\lambda=1$, and $g_{0}=0$, we estimate this threshold is $\Delta v^{*}=0.08(2)$.
@@
Finally, it is worth emphasizing that the derived stochastic continuum equations, Eq. (III), contain just the minimal set of drive-dependent couplings, which, at the mean field level, cannot produce the core phenomenology of the lattice-based models such as the discontinuities and anisotropies of the structure factors. Such phenomena would normally only be accessible through renormalization group analysis of the structure factor corrections due to the noise terms and nonlinearities. Nevertheless, by (numerically) seeking steady-state solutions of the stochastic equations, we can access these corrections and find behavior analogous to the nonequilibrium lattice models. Due to both analytical and computational limitations, our study has focused primarily on qualitative, comparative analyses. Nevertheless, we hope that this work provides a useful starting point for developing a more complete field-theoretical framework for this broader class of nonequilibrium systems.
@@ Acknowledgements.
Appendix A Gradient expansion
@@
where $\Delta_{t}$ stands for the discrete time-step taken, and $\Delta_{\mathrm{NN}}J_{A}^{\mathrm{NN}}$ for the discrete divergence associated with NN hops, etc. Explicitly, the $\Delta_{\mathrm{NN}}J$’s (with $t$ suppressed) are:
$-\Delta_{\mathrm{NN}}J_{A}^{\mathrm{NN}}\left(i,j\right)$ $=$ $\frac{1}{8}\sum_{\kappa=\pm 1}\bigg{\{}\rho_{A}(i-\kappa,j)h(i,j)\not{\rho}_{B}(i,j;\kappa)-\rho_{A}(i,j)h(i+\kappa,j)\not{\rho}_{B}(i+\kappa,j;\kappa)\bigg{\}}+i\leftrightarrow j$
$-\Delta_{\mathrm{NN}}J_{A}^{\delta,\mathrm{NN}}(i,j)$ $=$ $\frac{1}{8}\sum_{\kappa=\pm 1}\kappa\bigg{\{}\rho_{A}(i-\kappa,j)h(i,j)\not{\rho}_{B}(i,j;\kappa)-\rho_{A}(i,j)h(i+\kappa,j)\not{\rho}_{B}(i+\kappa,j;\kappa)\bigg{\}}$ (23)
@@
$\not{\rho}_{B}(i,j;\kappa)=(1$ $-\rho_{B}(i+\kappa,j))$
$(1-\rho_{B}(i,j+\kappa))(1-\rho_{B}(i,j-\kappa)).$
accounts for the NN AA-$B$ exclusion. There are analogous expressions for $\Delta_{\mathrm{NNN}}J$’s; see [48] for details. From these, we can construct evolution equations for the deviations (from $\rho$ and 0) of the density and charge fields (defined in Eq. (5)) $\phi_{\pm}\left(\mathbf{s},t\right)$. Let us write these symbolically as
@@
Since the $J$’s are polynomials in $\rho_{A,B}$ (up to sixth order), expressions like those in Eq. (A) are polynomial up to $O\left(\phi^{6}\right)$. Further, the right hand side of Eq. (24) is a divergence, so the coefficients in the polynomials may range up to $O\left(\rho^{5}\right)$. Needless to say, writing out all the orders for the resulting expressions for the right hand side is a tedious task, albeit a straightforward one. To arrive at the more familiar form of Eq. (13), we must take the continuum limit and, to simplify the divergence terms, keep only the lowest terms in the gradients. The remainder of this Appendix provides one possible approach.
Consider our lattice to have spacing $\ell$ in continuous space ($\mathbf{r}=\left(x,y\right)$), so that the fields above are defined on $\mathbf{r}=\mathbf{s}\ell$. One way to find a continuum limit is to let $\ell\rightarrow 0$ and assume that $\phi_{\pm}\left(x,y\right)$ is a smooth function, so that, e.g.,
@@
Combining both sides and keeping the lowest orders in both the fields and gradients, we obtain a set of equations similar to the deterministic parts of Eq. (III):
@@
$\qquad+\lambda_{+}\nabla^{2}\phi_{+}^{2}+\lambda_{-}\nabla^{2}\phi_{-}^{2}+g_{+}\partial_{x}\phi_{+}^{2}+g_{-}\partial_{x}\phi_{-}^{2},$
@@
$\qquad+\tilde{\lambda}_{+}\nabla(\phi_{-}\nabla\phi_{+})+\tilde{\lambda}_{-}\nabla(\phi_{+}\nabla\phi_{-})+g_{0}\partial_{x}(\phi_{+}\phi_{-})$

### v2

@@ V Discussion and conclusion
We investigate a model of a driven binary mixture (dubbed the “FLM”) which effectively captures the behavior of the driven Widom-Rowlinson lattice gas in both the low- and high-density phases, but includes an additional intermediate regime with irregular stripes. In the high-density, ordered phase, earlier studies [1, 2] report standard coarsening dynamics in the absence of a drive, while the drive leads to the emergence of regular stripes oriented perpendicular to the drive. In Sec. II, we briefly characterized the FLM and observed striking similarities with the DWRLG phenomenology, suggesting that both models may belong to the same universality class within a properly chosen parameter space. Notably, we observe that perpendicular stripes emerge even when the system is initialized with parallel stripes [43], indicating a robust stripe-selection mechanism. This reorientation appears to proceed via an instability that is reminiscent of the Mullins-Sekerka instability of the interfaces [43].
@@
Above the critical density, roughly estimated as $\rho_{c}\simeq 0.39$ for $\delta=0$ and $\rho_{\ell}=0.40$ for $\delta=1$, we identify two additional transition densities in the FLM: (i) For $\delta>0$, the density $\rho_{u}$ ($\simeq 0.65$, for $\delta=1$) marks a transition from short-range ordered (irregular) stripes to long-range ordered (regular) stripes [see, e.g., Figs. 1(a,d)- 5(b,c)]. The irregular stripe region, $\rho_{\ell}<\rho<\rho_{u}$, is characterized by pronounced long-wavelength instabilities, which serve as a mechanism to reduce the number of stripes. (ii) For $\delta\geq 0$, $\rho_{g}\simeq 0.90$; marks the onset of a frozen regime for uniform initial conditions [Figs. 1(a,b)], where coarsening arrests and glassy configurations appear. In Sec. II we offer a number of ways to characterize these two regimes.
@@
In the driven low-density regime, the FLM shows indications of a peak in the CF structure factor [upper panel in Fig. 6(a)] at an $L$-independent characteristic wavevector, $q^{*}=0.101\pm 0.001$. The DF structure factor [lower panel in Fig. 6(a)] also shows a shoulder at $2q^{*}_{-}$, but remains peaked at $2\pi/L$, consistent with the results reported in Refs. [1, 2]. Examining a snapshot of a typical configuration in this regime [Fig. 6(b)], we see that this microemulsion phase emerges from a tendency to form stripes perpendicular to the drive that is frustrated by the stronger effect of the drive in regions having low particle density.
@@
We derive a continuum version of the FLM via a gradient expansion in Sec. III. The analysis yields a system of two (reciprocally) coupled nonlinear stochastic partial differential equations, Eq. (III). These equations are consistent with the field equations developed in Ref. [2] for the low-density phase, but include additional terms that are important for describing the high-density regime, namely the terms proportional to $\lambda$ and $\Gamma_{\pm}$ in Eq. (III). Additionally, the charge fluctuation diffusion constant, $D_{-}$, changes sign at $\rho_{c}^{\mathrm{FT}}=0.37\dots$, marking the onset of the transition from a disordered to an ordered regime. For comparison, the FLM estimates are $\rho_{c}^{(0)}\simeq 0.39$ for the critical density at zero drive ($\delta=0$) and $\rho_{\ell}\simeq 0.40$ as the critical density for (irregular) stripe formation at $\delta=1$.
@@
In Sec. IV, we integrate the field equations numerically. We find that the PSSETD2H method is best suited for this task  [^16] Interestingly, we found no references in literature that describe these methods in detail, leading us to believe that such a description (and application) could be useful, particularly in the active matter context, since the non-reciprocal Cahn-Hilliard and Active Model B equations are analogous to Eq. (III). Dealiasing techniques are straightforward to implement, effectively minimizing errors from spatial discretization, which is particularly important in the presence of noise  [^17] In fact, we find that it is necessary to use a much finer mesh in space and time than the one we use in Sec. IV to numerically integrate an equation of type (III) with a noise of moderate intensity, $\sigma\approx 10^{-2}$, without recurring to dealiasing techniques.. The remainder of Sec. IV provides a qualitative overview of the continuum stochastic field theory, with minimal emphasis on quantitative analysis. A central aspect is that noise plays a crucial role in reproducing the features observed in previous studies. We observe not only a discontinuity in the structure factors $S_{\pm}(\mathbf{q})$ at the origin but also a peak in the charge fluctuation structure factor $S_{-}(\mathbf{q})$ at a nonzero wavevector along the drive direction, an indicator of the microemulsion phase.
@@
Turning to the high-density regime, the undriven ($\delta=0$) case displays standard coarsening dynamics consistent with the Lifshitz-Slyozov law in both the lattice model and the field theory. Our main findings for $\delta>0$ are summarized in Figs. 10-11: We identify a regime in the field theory, in which the equations of motion best reproduce the features of the lattice models (see Fig. 11 (2)). In this regime, as both density and drive increase, the continuum model transitions from irregular- to regular-stripe configurations, similarly to the FLM. Additionally, we uncover several novel regimes not previously reported in the lattice studies, including stripes aligned parallel to the drive and turbulent-like behavior (see Fig. 13).
@@
A key finding of Sec. IV is that a nonzero difference in the characteristic velocities of the fields, $\Delta v$, is a necessary condition for the formation of stripes aligned perpendicular to the drive, analogously to what is found for the microemulsion phase behavior in Ref. [2]. When $\Delta v=0$, such stripes do not form, even when the other drive-dependent couplings $g_{0}$ and $g_{-}$ are nonzero. Moreover, the modulation of the CF field (and, consequently, the DF field) along the drive direction, $q^{*}$, increases with the magnitude of $\Delta v$, in agreement with findings from the low-density regime in [2]; see Fig. 12. However, vertically aligned stripes emerge only when $\Delta v$ exceeds a finite threshold.
@@
Finally, it is worth emphasizing that the stochastic continuum equations, Eq. (III), contain just the minimal set of drive-dependent couplings, which, at the mean field level, cannot produce the core phenomenology of the lattice-based models such as the discontinuities and anisotropies of the structure factors. Such phenomena would normally only be accessible through renormalization group analysis of the structure factor corrections due to the noise terms and nonlinearities. Nevertheless, by (numerically) seeking steady-state solutions of the stochastic equations, we effectively access these corrections and find behavior analogous to the nonequilibrium lattice models. Due to both analytical and computational limitations, our study has focused primarily on qualitative, comparative analyses. Nevertheless, we hope that this work provides a useful starting point for developing a more complete field-theoretical framework for this broader class of nonequilibrium systems.
@@ Acknowledgements.
Appendix A Gradient expansion I – equations of motion
@@
$\Delta_{t}\rho_{A}(\mathbf{s},t)\equiv\rho_{A}(\mathbf{s},t+\tau)-\rho_{A}(\mathbf{s},t)$
is the change in the field at site $\mathbf{s}$ over a time increment $\tau$, and $\Delta_{\mathrm{NN}}J_{A}^{\mathrm{NN}}$ is the discrete divergence associated with NN hops, etc. Explicitly, the $\Delta_{\mathrm{NN}}J$’s (with $t$ suppressed) are:
$-\Delta_{\mathrm{NN}}J_{A}^{\mathrm{NN}}\left(i,j\right)$ $=\frac{1}{8}\sum_{\kappa=\pm 1}\bigg\{\rho_{A}(i-\kappa,j)h(i,j)\not{\rho}^{\leftrightarrow}_{B}(i,j;\kappa)-\rho_{A}(i,j)h(i+\kappa,j)\not{\rho}^{\leftrightarrow}_{B}(i+\kappa,j;\kappa)\bigg\}$
${}\ \ \ +\frac{1}{8}\sum_{\kappa=\pm 1}\bigg\{\rho_{A}(i,j-\kappa)h(i,j)\not{\rho}^{\updownarrow}_{B}(i,j;\kappa)-\rho_{A}(i,j)h(i,j+\kappa)\not{\rho}^{\updownarrow}_{B}(i,j+\kappa;\kappa)\bigg\},$
$-\Delta_{\mathrm{NN}}J_{A}^{\delta,\mathrm{NN}}(i,j)$ $=\frac{1}{8}\sum_{\kappa=\pm 1}\kappa\bigg\{\rho_{A}(i-\kappa,j)h(i,j)\not{\rho}^{\leftrightarrow}_{B}(i,j;\kappa)-\rho_{A}(i,j)h(i+\kappa,j)\not{\rho}^{\leftrightarrow}_{B}(i+\kappa,j;\kappa)\bigg\},$ (19)
where
$\not{\rho}^{\leftrightarrow}_{B}(i,j;\kappa)=(1$ $-\rho_{B}(i+\kappa,j))$
$(1-\rho_{B}(i,j+\kappa))(1-\rho_{B}(i,j-\kappa)),$
$\not{\rho}^{\updownarrow}_{B}(i,j;\kappa)=(1$ $-\rho_{B}(i,j+\kappa))$
$(1-\rho_{B}(i+\kappa,j))(1-\rho_{B}(i-\kappa,j))$
account for the NN AA-$B$ exclusion for horizontal ($\leftrightarrow$) and vertical ($\updownarrow$) jumps, respectively, and $h(i,j)\equiv 1-\rho_{A}(i,j)-\rho_{B}(i,j)$ for the density of holes at site $\mathbf{s}=(i,j)$. There are analogous expressions for $\Delta_{\mathrm{NNN}}J$’s; see [43] for details. From these, we can construct evolution equations for the deviations (from $\rho$ and 0) of the density and charge fields (defined in Eq. (5)) $\phi_{\pm}\left(\mathbf{s},t\right)$. Let us write these symbolically as
@@
Since the $J$’s are polynomials in $\rho_{A,B}$ (up to sixth order), expressions like those in Eq. (A) are polynomial up to $O\left(\phi^{6}\right)$. To arrive at the more familiar form of Eq. (13), we must take the continuum limit and, to simplify the divergence terms, keep only the terms of lowest order in the gradients. The remainder of this Appendix provides one possible approach.
Consider our lattice to have spacing $\ell$ in continuous space ($\mathbf{r}=\left(x,y\right)$), so that the fields above are defined on $\mathbf{r}=\mathbf{s}\ell$. One path to a continuum limit is to let $\ell\rightarrow 0$ and assume that $\phi_{\pm}\left(x,y\right)$ is a smooth function, so that, e.g.,
@@
Combining both sides and keeping only the relevant terms to fourth order in the derivatives and to second order in both fields, we obtain a set of equations similar to the deterministic parts of Eq. (III):
@@
$\hskip 18.49988pt+\lambda_{+}\nabla^{2}\phi_{+}^{2}+\lambda_{-}\nabla^{2}\phi_{-}^{2}+g_{+}\partial_{x}\phi_{+}^{2}+g_{-}\partial_{x}\phi_{-}^{2},$
@@
$\hskip 18.49988pt+\tilde{\lambda}_{+}\nabla(\phi_{-}\nabla\phi_{+})+\tilde{\lambda}_{-}\nabla(\phi_{+}\nabla\phi_{-})+g_{0}\partial_{x}(\phi_{+}\phi_{-})$
From the above equation, we see that the undriven currents appearing in Eq. (13), $\mathbf{J}_{\pm}$, read
while the drive-dependent currents, $J_{\pm}^{\delta}$, are given by
$J^{\delta}_{+}=-v_{+}\phi_{+}+g_{+}\phi_{+}^{2}+g_{-}\phi_{-}^{2}$
$J^{\delta}_{-}=-v_{-}\phi_{-}+g_{0}\phi_{+}\phi_{-}.$ (22)
Note that the form of the terms in Eqs. (A)-(A) is constrained by the $A\leftrightarrow B$ symmetry which implies that $\mathbf{J}_{+}$ ($\mathbf{J}_{-}$) cannot have odd (even) powers of $\phi_{-}$. Let us briefly highlight some characteristics of qualitative importance. First, based on this continuum limit of the FLM, we find that $D_{+}$ is positive for all $\rho$, which implies we can expect the DF to be stable, unless it is driven unstable through coupling to $\phi_{-}$. On the other hand, as expected of systems that display coarsening dynamics, $D_{-}$ changes sign when $\rho$ exceeds a critical value.
Exploiting a Galilean transform, we will work in the co-moving frame of the DF, so that the $v_{+}\phi_{+}$ term is absent in $J_{+}^{\delta}$ and $v_{-}\phi_{-}$ is replaced by $(\Delta v)\phi_{-}$, where $\Delta v\equiv v_{-}-v_{+}$. Note that both currents $J_{\pm}^{\delta}$ vanish when $\delta=0$, along with all their individual $v$ and $g$ terms.
Finally, we must add noise terms, $\xi_{\pm}$, which are difficult to derive from first principles, but we know they must conserve the total particle number and have zero mean. Taking the simplest possible approach, let $\xi_{A}$ and $\xi_{B}$ be the noise terms in the equations for $\rho_{A}$ and $\rho_{B}$. To first approximation, these are independent zero-mean conserved Gaussian noises, with variances $\sigma_{A}^{2}$ and $\sigma_{B}^{2}$. By symmetry under exchange of species $A$ and $B$, we must have $\sigma_{A}=\sigma_{B}$. In this case,
$\xi_{+}=\xi_{A}+\xi_{B},\qquad\xi_{-}=\xi_{A}-\xi_{B},$ (23)
have the same variance $2\sigma^{2}_{A}$, and are uncorrelated, $\langle\xi_{+}\xi_{-}\rangle=0$. Thus, it follows that $\xi_{\pm}$ are zero-mean conserved Gaussian noise terms, uncorrelated in time and with each other, having variance $\sigma^{2}\equiv 2\sigma^{2}_{A}$, i.e.,
$\langle\xi_{\pm}(\mathbf{r},t)\rangle=0$
$\langle\xi_{+}(\mathbf{r},t)\xi_{-}(\mathbf{r}^{\prime},t^{\prime})\rangle=0$
$\langle\xi_{\pm}(\mathbf{r},t)\xi_{\pm}(\mathbf{r}^{\prime},t^{\prime})\rangle=-\sigma^{2}\nabla^{2}\delta(\mathbf{r}-\mathbf{r}^{\prime})\delta(t-t^{\prime}).$ (24)
Combining all the ingredients, we arrive at the SPDEs
$\partial_{t}\phi_{+}$ $=D_{+}\nabla^{2}\phi_{+}-\Gamma_{+}\nabla^{4}\phi_{+}$
$+\lambda_{+}\nabla^{2}\phi_{+}^{2}+\lambda_{-}\nabla^{2}\phi_{-}^{2}+g_{+}\partial_{x}\phi_{+}^{2}+g_{-}\partial_{x}\phi_{-}^{2}+\xi_{+},$
$+\tilde{\lambda}_{+}\nabla(\phi_{-}\nabla\phi_{+})+\tilde{\lambda}_{-}\nabla(\phi_{+}\nabla\phi_{-})+g_{0}\partial_{x}(\phi_{+}\phi_{-})+\xi_{-}$ (25)
For convenience, we collected all the linear/quadratic terms on the first/second line. In Appendix B, we show in some detail how to calculate the exact expressions of the linear coefficients. We also show how a set of assumptions is used to simplify the nonlinear terms in Eq. (25) to arrive at Eq. (III).
Appendix B Gradient expansion II – coefficients
To calculate the exact expression of the linear coefficients, we start by substituting
$2\rho_{A}=\phi_{+}+\phi_{-}+\rho$
$2\rho_{B}=\phi_{+}-\phi_{-}+\rho$
in Eq. (18) to write the expressions in Eq. (A) (and the analogous expressions for species B) in terms of the density and charge fluctuations, but this time we only keep the resulting linear piece. Because the change of variables above involves the parameter $\rho$ (which is zeroth order in the fields), and since the $J$’s are sixth-order polynomials in $\rho_{A,B}$, the linear terms will have contributions coming from all possible orders in the fields. As a result, the corresponding linear coefficients are fifth-order polynomials in $\rho$.
The lattice update appearing in the different terms of Eq. (A) can be written in Fourier space in terms of the operator $e^{i\kappa\left(q_{x}+q_{y}\right)}$, such that for any field $f$ on the lattice,
$\mathcal{F}^{-1}\{e^{i\kappa\left(q_{x}+q_{y}\right)}\widetilde{f}(q_{x},q_{y})\}=f(i+\kappa,j+\kappa),$ (26)
where $\mathcal{F}^{-1}$ is the inverse discrete Fourier transform.
We take the discrete Fourier transform of the linear part of Eq. (A) and use the above property to express the lattice-updated fields using the exponentials. There will be analogous expressions for NNN hops. We keep working in Fourier space to substitute these expressions in Eq. (18) and write
$\Delta_{t}\phi_{\pm}(\mathbf{q},t)=\mathcal{L}_{\pm}(\mathbf{q})\phi_{\pm}(\mathbf{q},t),$
where
$\mathcal{L}_{\pm}(\mathbf{q})=$ $A_{\pm}+B_{\pm}\big(\cos\left(q_{x}\right)+\cos\left(q_{y}\right)\big)+C_{\pm}\big(\cos^{2}\left(q_{x}\right)+\cos^{2}\left(q_{y}\right)\big)+E_{\pm}\cos\left(q_{x}\right)\cos\left(q_{y}\right)$
$+G_{\pm}\big(\cos\left(q_{x}\right)+\cos\left(q_{y}\right)\big)\cos(q_{x})\cos(q_{y})$
$+i\delta\big(H_{\pm}+d_{1}\cos\left(q_{x}\right)+K_{\pm}\cos\left(q_{y}\right)+d_{2}\cos\left(q_{x}\right)\cos\left(q_{y}\right)+d_{3}\cos^{2}\left(q_{y}\right)\big)\sin\left(q_{x}\right).$ (27)
All coefficients are polynomials in $\rho$; their explicit forms are omitted for clarity and conciseness. Up to this point, no approximations have been introduced in the linear part: these are exact expressions, with $\mathcal{L}_{\pm}(\mathbf{q})$ denoting the exact linear operators corresponding to the Fourier-space representation of the linear terms in the mass-transfer equation (18). The gradient expansion (GE) of Eq. (27) yields, to lowest order,
$\mathcal{L}_{\pm}(\mathbf{q})\approx-D_{\pm}|\mathbf{q}|^{2}-\Gamma_{\pm}|\mathbf{q}|^{4}-v_{\pm}\,iq_{x}.$ (28)
Comparing with expression (27) and collecting similar terms, we find the following coefficients  [^18] For simplicity, we suppressed all the “physical units” for the coefficients, e.g., $D\propto\ell^{2}\tau^{-1}$, etc. Their detailed forms play no role in the numerical studies.:
$D_{+}$ $=\frac{2}{64}(2-\rho)^{2}\left(3+4\rho+2\rho^{3}-8\rho^{2}\right),$
$v_{+}$ $=-\frac{\delta}{64}(2-\rho)^{2}\left(-4+26\rho-29\rho^{2}+8\rho^{3}\right),$
$v_{-}$ $=\frac{3\,\delta}{32}(2-\rho)^{2}\left(2-3\rho^{2}+\rho^{3}\right),$
$\Delta v$ $\equiv v_{-}-v_{+}=\frac{\delta}{64}(2-\rho)^{2}\left(8+26\rho-47\rho^{2}+14\rho^{3}\right).$ (29)